OpenStudy (anonymous):

Find the midpoint of the segment joining the points: (4, -2) and (-8, 6)

OpenStudy (ipwnbunnies):

The midpoint of two points is simply: \[(\frac{x1+x2}{2} , \frac{y1+y2}{2})\]

OpenStudy (ipwnbunnies):

Where (x1,y1) and (x2,y2) are the two points.

OpenStudy (anonymous):

\[ \left( \frac{ x _{1}+x _{2} }{ 2 },\frac{ y _{1}+y _{2} }{ 2 }\right)\] You can use this with the little subscripts

OpenStudy (anonymous):

Next time because I'm a subscript nazi

OpenStudy (ipwnbunnies):

LOL.